A ball thrown into the air reaches a height h (in feet) according to the formula h = -16t2 + 80t, where t is the time elapsed since the ball was thrown. Find the shortest time it will take the ball to reach a height of 64 feet.

Answer by John10(297)

when the height reaches to 64 ft,
-16t^2 + 80t = 64
-16t^2 + 80t - 64 = 0
t^2 - 5t + 4 = 0
(t - 4) (t - 1) = 0
t = 4 or t = 1
thus the shortest time is t = 1 second
John10:)